# Online Resource Allocation via Static Bundle Pricing

### v1

@@ Abstract
Online Resource Allocation addresses the problem of efficiently allocating limited resources to buyers with incomplete knowledge of future requests. In our setting, buyers arrive sequentially demanding a set of items, each with a value drawn from a known distribution. We study environments where buyers’ valuations exhibit complementarities. In such settings, standard item-pricing mechanisms fail to leverage item multiplicities, while existing static bundle-pricing mechanisms rely on problem-specific arguments that do not generalize.
We develop a unified technique for online resource allocation with complementarities for three domains: (i) single-minded combinatorial auctions with maximum bundle size $d$, (ii) general single-minded combinatorial auctions, and (iii) a graph-based routing model in which buyers request to route a unit of flow from a source node $s$ to a target node $t$ in a capacitated graph. Our approach yields static and anonymous bundle-pricing mechanisms whose performance improves exponentially with item capacities. For the $d$-single-minded setting with minimum item capacity $B$, we obtain an $O(d^{1/B})$-competitive mechanism, recovering the known $O(d)$ bound for unit capacities ($B=1$) and achieving exponentially better guarantees as capacities grow. For general single-minded combinatorial auctions and the graph-routing model, we obtain $O(m^{1/(B+1)})$-competitive mechanisms, where $m$ is the number of items.
We complement these results with information-theoretic lower bounds. We show that no online algorithm can achieve a competitive ratio better than $\Omega((m/\ln m)^{1/(B+2)})$ in the general single-minded setting and $\Omega((d/\ln d)^{1/(B+1)})$ in the $d$-single-minded setting. In doing so, we reveal a deep connection to the classical extremal combinatorics problem of determining the maximum number of qualitatively independent partitions of a ground set.
footnotetext: This work has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0, funded by the European Union under the NextGenerationEU Program.
@@ 1 Introduction
Online Resource Allocation is a central problem in Economics and Operations Research. A seller faces a sequence of buyers, each demanding a set of resources. The seller must make irrevocable allocation decisions with limited information about future arrivals. In this work, we study the setting in which the seller has access to the distribution of buyers’ valuations, placing the problem within the prophet-inequality framework.
@@
The study of prophet inequalities originates in the classical single-item setting. Krengel, Sucheston, and Garling [krengel1977] and independently Samuel-Cahn [samuel1984comparison] established the foundational results comparing online stopping rules to a prophet who observes all realizations in advance. Hajiaghayi et al. [haji1007] later connected prophet inequalities to mechanism design, showing how prophet-style guarantees can be leveraged to design truthful posted-price mechanisms. Building on this connection, Chawla et al. [chawla2009] developed a theory of sequential posted pricing mechanisms.
@@
Following these foundational results, research on prophet inequalities expanded rapidly, branching into richer allocation environments and more complex valuation models. In particular, combinatorial auctions emerged as a central and intensively studied setting within the prophet-inequality framework. For submodular and XOS valuations, Feldman et al. [fgl-balanced14] gave a $2$-competitive posted-price mechanism. For subadditive valuations, Dütting et al. [duttingloglogm] obtained, via item-pricing, an $O(\log\log m)$-competitive mechanism (where $m$ is the number of items), a guarantee that was recently improved by Correa and Cristi [Correa2023ACF], who achieved a $6$-competitive prophet inequality. For MPH-$k$ valuations, Braun and Kesselheim [braun2022simplifiedprophetinequalitiescombinatorial] designed an item-pricing mechanism with competitive ratio $2k+\sqrt{2k(k-1)}-1$.
@@
In this work, we focus on single-minded combinatorial auctions, a setting that captures strong complementarities: each buyer derives value only if they obtain their entire desired bundle. Single-minded combinatorial auctions capture the most extreme form of complementarities and arise naturally in applications where partial allocation provides no utility. Their simplicity makes them a fundamental testbed for understanding how complementarities interact with online decision-making and posted-price mechanisms. A canonical example is pricing goods in a grocery store. A shopper attempting to bake a cake gains value only if they obtain all necessary ingredients; eggs and milk provide zero utility if flour is missing.
@@
Extending the grocery store analogy, many real systems feature multiple copies of each item rather than unique goods. A growing line of work examines how such multiplicities affect the performance of pricing mechanisms. Alaei [multi-alei] showed that, in the single-item setting with capacity $k$, one can achieve $1-O(1/\sqrt{k})$ competitiveness. More recent results by Jiang et al. [multi-jiang] and Chawla et al. [multi-suchi] give matching upper and lower bounds for all values of $k$, while Chawla et al. [chawla25] continued this line of work for XOS valuations. Recently, Ghuge et al. [ghuge2025] studied online resource allocation with large capacities under stochastic arrivals using dynamic pricing mechanisms. They show that near-optimal welfare can be achieved using only a single sample from each distribution, highlighting that high capacities substantially mitigate uncertainty even with limited information.
@@
In contrast, our work focuses on static mechanisms, in which the seller commits to a menu of prices before any buyer arrives and does not adjust them over time. Although such mechanisms are simple, transparent, and widely deployed in practice, they face significant limitations in the presence of complementarities. The seminal work of Feldman et al. [fgl-balanced14] established strong impossibility results for item pricing in combinatorial auctions: when buyers’ valuations require complements, no item-pricing mechanism can be less than $\Omega(m)$-competitive. When parameterizing by the maximum bundle size $d$, Correa et al. [correa-d] obtained a tight $(d+1)$-competitive guarantee.
Motivated by these challenges, Chawla et al. [chawla18] initiated the study of bundle pricing in settings with complementarities, first in the context of online allocation of intervals and later for paths in trees. Their results demonstrated that richer pricing menus can circumvent the barriers inherent to item pricing. Moreover, their work highlighted that static pricing mechanisms admit progressively stronger guarantees as the number of available copies of each resource increases.
This raises two fundamental questions:
What static pricing mechanisms are appropriate for environments with complementarities?
How does the availability of multiple item copies affect the achievable performance of such mechanisms?
In this work, we develop a unified approach for prophet inequalities in the presence of complementarities. Our mechanisms are static (the corresponding menus are fully determined before any arrivals occur) and anonymous (buyers face the same menu). We show that this approach is applicable to the following domains:
@@
    the $d$-single-minded setting, where each buyer demands a bundle of size at most $d$
@@
    general single-minded combinatorial auctions, with arbitrary bundle sizes
@@
    the graph-routing setting, where buyers seek any feasible $s$-$t$ path in a capacitated graph.
Across all of these environments, a pattern emerges: bundle prices become exponentially more powerful as item capacities increase. Even small increases in capacity can dramatically improve achievable competitive ratios, revealing a sharp structural transition in the difficulty of online allocation with complementarities.
At the same time, we complement our positive results with information-theoretic lower bounds for both the $d$-single-minded and the general single-minded settings. Our lower-bound construction is obtained by embedding the classical hard instance of Babaioff et al. [babaioff] into the single-minded combinatorial auction framework. This embedding reveals an intriguing connection between our impossibility results and an extremal combinatorics problem: determining the maximum number of pairwise qualitatively independent partitions of a ground set, a question first studied by Rényi [renyi]. Together, these lower bounds show that the dependence of our guarantees on the number of items $m$ and the minimum multiplicity $B$ is essentially unavoidable.
1.1 Technical Overview and Contributions
We study online allocation problems with strong complementarities under distributional information, focusing on static and anonymous bundle-pricing mechanisms. All guarantees hold in expectation over both the randomness of the mechanism and buyers’ realized values, and the arrival order is chosen by an almighty adversary who observes the entire realization and the mechanism’s internal randomness.
Our starting point is the ex-ante linear relaxation, which upper bounds the expected welfare of the prophet, i.e. the expected value over the realizations welfare of the optimal allocation. Before solving this LP, we scale down all item capacities by a factor $\gamma$, creating a slack that later absorbs the fluctuations created by the randomness of the realization. This scaled relaxation guides the design of our pricing mechanism.
A notable feature of the optimal fractional solution is its simple and rigid structure. For any buyer, the LP allocates probability mass to values in an almost threshold-like form: below some cutoff the allocation is zero, above it the LP fully saturates the value distribution, and at most one crucial value may receive partial mass. When aggregated over all buyers who demand the same bundle $S$, these patterns define a canonical price level $w_{S}$ that dictates how the menu treats $S$.
This structure directly informs our menu construction. The mechanism posts a static, anonymous menu containing multiple copies of each bundle $S$, priced either at the crucial value $w_{S}$ or at the next higher value in the buyers’ value support, $w_{S}+1$. The guiding objective is to emulate the fractional solution as closely as possible: ideally, a buyer $b$ with realized value $v$ would find an available copy of their desired bundle with probability $x_{b,v}/q_{b,v}$, independently of other buyers, where $x_{b,v}$ is the LP mass assigned to the buyer-value pair $(b,v)$ and $q_{b,v}=\mathbb{P}\!\left[v_{b}=v\right]$.
Such independent acceptance probabilities, however, cannot be implemented directly using a static and anonymous menu. Instead, the mechanism must commit ex-ante to a fixed number of copies of each bundle at each price. Guided by the LP solution, we post no copies of bundle $S$ at prices below $w_{S}$, and essentially unlimited copies at prices strictly above $w_{S}$, relying on the scarcity of high-value buyers to limit congestion. The number of copies offered exactly at price $w_{S}$ is more delicate. Offering too many copies risks exhausting item capacities and blocking high-value buyers later in the sequence, while offering too few results in a significant loss relative to the fractional benchmark. To balance these effects, we randomize between offering $\lfloor x_{S,w_{S}}\rfloor$ and $\lceil x_{S,w_{S}}\rceil$ copies of bundle $S$ at price $w_{S}$, where $x_{S,w_{S}}$ denotes the total LP mass assigned to buyers demanding $S$ with value $w_{S}$.
A desirable property of our menu is price subadditivity: for every bundle $S$, the posted price of $S$ should be no greater than the total price of any collection of bundles whose union contains $S$. This guarantees that a buyer who desires $S$ always prefers purchasing a copy of $S$ rather than assembling a larger covering collection. Enforcing price subadditivity, however, is nontrivial. The prices suggested by the LP solution need not satisfy this property, and the buyers’ original value supports may not contain appropriate price levels that would permit a subadditive menu without sacrificing performance. To address this issue, we first augment each buyer’s value support with carefully chosen auxiliary price points. We then apply a post-processing step to the LP solution that enforces subadditivity before executing the menu-construction procedure described above.
To analyze the mechanism, we introduce an unconstrained algorithm: a hypothetical process that accepts every buyer who selects an available menu entry, ignoring item-capacity constraints. This auxiliary algorithm is easier to analyze and satisfies two key properties:
@@ 1.1 Technical Overview and Contributions
    Its expected welfare is a constant fraction of the scaled LP optimum, denoted by $\textnormal{{FracOpt}}_{\gamma}$ throughout the paper.
@@
    For every item $e$, the expected number of sold copies that use $e$ is at most $c(e)/\gamma$.
The welfare of the actual mechanism equals the welfare of the unconstrained algorithm minus the value of blocked buyers-those accepted by the unconstrained algorithm but rejected when enforcing capacity constraints. Property (1) implies that it suffices to bound the expected welfare lost due to blocking as a constant fraction of the unconstrained value, while Property (2) allows us to control the expected contribution of blocked buyers.

### v2

@@ Abstract
Online Resource Allocation addresses the problem of efficiently allocating limited resources to buyers with incomplete knowledge of future requests. In our setting, buyers arrive sequentially requesting a set of items, each with a value drawn from a known distribution. We study the efficiency of static and anonymous bundle pricing in environments where the buyers’ valuations exhibit strong complementarities. In such settings, standard item pricing fails to leverage item multiplicities, while static bundle pricing mechanisms are only known for very restricted domains and their analysis relies on domain-specific arguments.
We develop a unified bundle pricing framework for online resource allocation in three well-studied domains with complementarities: (i) single-minded combinatorial auctions with maximum bundle size $d$; (ii) general single-minded combinatorial auctions; and (iii) network routing, where each buyer aims to route a unit of flow from a source node $s$ to a target node $t$ in a capacitated network. Our approach yields static and anonymous bundle pricing mechanisms whose performance improves exponentially with item multiplicity. For the $d$-single-minded setting with minimum item multiplicity $B$, we obtain an $O(d^{1/B})$-competitive mechanism. For general single-minded combinatorial auctions and online network routing, we obtain $O(m^{1/(B+1)})$-competitive mechanisms, where $m$ is the number of items.
We complement these results with information-theoretic lower bounds. We show that no online algorithm can achieve a competitive ratio better than $\widetilde{\Omega}(m^{1/(B+2)})$ for single-minded combinatorial auctions and $\widetilde{\Omega}(d^{1/(B+1)})$ for the $d$-single-minded setting. Our constructions exploit a deep connection to the classical extremal combinatorics problem of determining the maximum number of qualitatively independent partitions of a ground set.
@@ 1 Introduction
Online Resource Allocation is a central problem in Economics and Operations Research. A seller faces a sequence of buyers, each demanding a set of resources, and must make irrevocable allocation decisions with limited information about future arrivals. In this work, we study online resource allocation in a Bayesian setting where the seller has access to the distribution of buyers’ valuations, placing the problem within the general prophet-inequality framework.
General Background.
@@
The study of prophet inequalities originates in the classical single-item setting. Krengel and Sucheston (1977) (also crediting D.J.H. Garling) and independently Samuel-Cahn (1984) established the foundational results comparing the value obtained by online stopping rules against the best possible (a.k.a. optimal) value obtained by a “prophet” who observes all realizations in advance. More recently Hajiaghayi et al. (2007) observed the connection of prophet inequalities to mechanism design, showing how prophet-style guarantees can be leveraged to design truthful posted-price mechanisms. Building on this connection, Chawla et al. (2010) initiated the study of prophet inequalities for revenue maximization, further expanding the framework’s applicability beyond the social welfare.
@@
Following these foundational results, research on prophet inequalities expanded rapidly, branching into richer allocation environments and more complex valuation models. In particular, combinatorial auctions emerged as a central and intensively studied setting within the prophet-inequality framework. For submodular and XOS valuations, Feldman et al. (2015) significantly generalized the basic prophet inequality, introducing the elegant framework of balanced prices, and obtained a $2$-competitive posted-item-price mechanism. For subadditive valuations, Dütting et al. (2024) obtained an $O(\log\log m)$-competitive item-pricing mechanism (where $m$ is the number of items), a guarantee recently improved by Correa and Cristi (2023), who achieved a $6$-competitive prophet inequality. For MPH-$d$ valuations, which exhibit a degree of complementarity quantified by $d$, Braun and Kesselheim (2022) presented an item-pricing mechanism with competitive ratio $2d+\sqrt{2d(d-1)}-1$.
@@
In this work, we focus on single-minded combinatorial auctions, which exhibit strong complementarities: buyers derive value only if they obtain their entire desired bundle. Single-minded combinatorial auctions capture the most extreme form of complementarities and arise naturally in many applications where partial allocation provides no value. Their simplicity makes them a fundamental testbed for understanding how complementarities interact with online decision-making and posted-price mechanisms. For a simple canonical example, we may refer to replenishing and pricing inventory in a grocery store. A shopper attempting to bake a cake gains value only if they obtain all necessary ingredients; eggs and milk provide zero utility if flour is missing.
@@
Extending the grocery store analogy, many practical applications feature multiple copies of each item rather than unique goods. A growing line of work examines how the performance of pricing mechanisms improves with item multiplicity. Alaei (2014) showed that in the single-item setting with capacity $B$, one can achieve $1+O(1/\sqrt{B})$ competitiveness. Subsequently, Jiang et al. (2023) and Chawla et al. (2023) gave matching upper and lower bounds for all values of $B$. Recently, Ghuge et al. (2025) studied online resource allocation with large capacities under stochastic arrivals using dynamic pricing mechanisms (see also e.g., Devanur et al. (2019); Chawla et al. (2017); Kesselheim et al. (2018); Gupta and Molinaro (2016) and their references for more previous work on achieving nearly optimal competitiveness for online resource allocation in the large-capacity regime). Ghuge et al. (2025) proved that for capacity $B=\widetilde{\Omega}(1/\varepsilon^{6})$, a competitive ratio of $1+O(\varepsilon)$ can be achieved using only a single sample from each distribution, highlighting that high capacities substantially mitigate uncertainty even with limited information.
Context and Motivation.
@@
We focus on static and anonymous pricing mechanisms, in which the seller commits to a menu of prices before any buyer arrives and does not adjust them over time. Although such mechanisms are simple, transparent and widely deployed in practice, they face significant limitations in the presence of complementarities. Feldman et al. (2015) proved that when buyers’ valuations exhibit complementarities, no item-pricing mechanism can be less than $\Omega(m)$-competitive. When parameterizing by the maximum bundle size $d$, Correa et al. (2023) obtained a tight $(d+1)$-competitive guarantee.
Motivated by these challenges, Chawla et al. (2019) initiated the study of bundle pricing in settings with strong complementarities, in the context of online allocation of intervals on the line and paths in trees. Their results demonstrated that richer pricing menus can circumvent the barriers inherent to item pricing. For allocating intervals, where the length of the longest interval requested is $d$ and items are available in $B$ copies, Chawla et al. (2019) gave an $O\left(\frac{\log d}{B\log\log d}\right)$-competitive static, anonymous bundle-pricing scheme and a matching lower bound on the performance of any online algorithm. For the more challenging setting of allocating paths in a tree, they presented a lower bound of $\Omega(\sqrt{d/\log d})$ on the competitiveness of any online algorithm and a static, anonymous bundle pricing scheme with competitive ratio $O((\log v_{\max})/B)$, where $v_{\max}$ is the maximum valuation. At a conceptual level, their results exploit the interval (resp. the tree) structure in order to reduce the setting with complementarities to the single item setting at a loss in the competitive ratio that scales with $\log d$ (resp. $\log v_{\max}$). Exploiting the restricted combinatorial structure of intervals, Chawla et al. (2019) improve exponentially on the best possible competitiveness of online allocations and show that the competitive ratio improves inversely with item capacities. Whether non-trivial competitive guarantees are possible for more general combinatorial structures (e.g., networks), even when the buyers have binary valuations and request bundles of equal size $d$, is posed as a challenging open question in Chawla et al. (2019).
Hence, previous work motivates the following fundamental questions:
@@
    What static pricing mechanisms best suit environments with complementarities?
@@
    How fast does the competitive ratio improve with item multiplicity?
@@
    Is there any standard framework for designing static bundle pricing mechanisms and establishing lower bounds in the presence of complementarities?
Contribution in a Nutshell.
In this work, we make substantial progress towards answering all the three questions above. On the positive side, we develop a unified framework for the design of bundle pricing mechanisms in the presence of complementarities. Our pricing mechanisms are static, in the sense that the corresponding pricing menus are fully determined before any buyer’s arrival, and anonymous, in the sense that the pricing menu presented to a buyer does not depend on their identity. We apply our framework to the following domains:
@@
    $d$-single-minded combinatorial auctions, where requested bundles are of size at most $d$.
@@
    General single-minded combinatorial auctions with arbitrary bundle sizes.
  3. 3.
    Online network routing, where each buyer $b$ aims to route a unit of flow from a source node $s_{b}$ to a destination node $t_{b}$ in an edge-capacitated network.
We obtain competitive ratios that are $O(d)$ and $O(\sqrt{m})$, respectively, in the basic case of unit item capacities, and improve exponentially with item capacities. Thus, one of our key conceptual findings is that bundle pricing becomes exponentially more powerful as item capacities increase (resembling what happens for the standard offline optimization versions of single-minded combinatorial auctions and multicommodity network routing, see e.g., Lehmann et al. (2002); Bartal et al. (2003); Kolliopoulos and Stein (2004); Lavi and Swamy (2011); Bansal et al. (2012)).
We complement our positive results with information-theoretic lower bounds for both the $d$-single-minded and the general single-minded settings that apply to any online algorithm. Our lower-bound construction is obtained by embedding the construction of Babaioff et al. (2007) into the setting of single-minded combinatorial auctions. Our proof builds on an intriguing connection to a fundamental question in extremal combinatorics: “what is the maximum number of pairwise qualitatively independent partitions of a ground set?”, first asked by Rényi (1970).
2 Model, Technical Overview and Main Results
Model.
We consider a set $\mathcal{M}$ of $m$ distinct items, each available in at least $B\in\mathbb{Z}_{>0}$ copies, and a set $\mathcal{N}$ of $n$ buyers, typically indexed by $b\in[n]$. For each buyer $b$, we are given a fixed set of items $S_{b}$ and a distribution $\mathcal{D}_{b}$ from which $b$’s value $v_{b}$ for their requested set $S_{b}$ is drawn. We assume that the distributions $\mathcal{D}_{b}$ are independent across buyers and are given to the mechanism in an explicit form, i.e., by the probabilities $q_{b,v}$ that buyer $b$ has value $v_{b}=v$ for getting their set $S_{b}$, for all buyers $b$ and all values $v$ in the support of $\mathcal{D}_{b}$ (which is assumed to be some fixed subset of non-negative integers). Each buyer $b$ has utility $v_{b}$ minus the price posted by the mechanism for their set $S_{b}$ (or a superset $S\supseteq S_{b})$, if $S_{b}$ (or some $S\supseteq S_{b}$) is allocated to $b$, and 0 otherwise.
We aim to maximize the expected social welfare, i.e., the expected total value of the buyers who receive their requested bundle, where expectation is taken over both the randomness of the mechanism and the buyers’ realized values. The mechanism’s competitive ratio is established against the expected social welfare of a “prophet”, who collects the expected (over the buyers’ realizations) total value of an optimal solution for the realized instance. Our mechanisms are online, in the sense that they have to irrevocably decide on the allocation of the requested bundles upon the agents’ arrival, without any knowledge of the realized values of the agents arriving in the future. The order in which the online mechanism deals with the buyers is determined by an almighty adversary that has access to the realized buyer values and the mechanism’s internal randomness.
2.1 General Framework for Static Bundle Pricing
Similar to previous work (see e.g., Chawla et al. (2019); Jiang et al. (2023); Lee and Singla (2018)), we start with the ex-ante linear programming relaxation (a.k.a. EA-LP, formally stated in Section 5), which upper bounds the expected social welfare of the prophet. Before solving EA-LP, we scale down all item capacities by a factor $\gamma$, creating a slack that later absorbs the demand fluctuations created by the randomness of the realization. The design of our pricing mechanism is guided by the optimal fractional solution to $\gamma$-scaled EA-LP.
A notable feature of EA-LP’s solution is its simple and rigid structure. For any buyer $b$, EA-LP allocates probability mass to values in an almost threshold-like form: below some cutoff value, the allocation is zero; above it, the allocation saturates the value distribution (resulting in so-called tight values); and there is at most one crucial value where the allocation is positive and non-tight (see Section 5.1). Aggregating over all buyers interested in the same bundle $S$, we define a canonical price level $w_{S}$ (a.k.a. the important value of $S$) which dictates how the pricing menu treats $S$.
This structure directly guides our menu construction (see Section 6). The mechanism posts a static, anonymous menu containing multiple “copies” of each bundle $S$, priced either at the important value $w_{S}$ or at the next higher value $w_{S}+1$. We aim to mimic the optimal fractional solution of EA-LP as closely as possible: ideally, a buyer $b$ with realized value $v$ should receive an available copy of their desired bundle with probability $x_{b,v}/q_{b,v}$, independently of other buyers, where $x_{b,v}$ is the allocation mass assigned by EA-LP to the buyer-value pair $(b,v)$ and $q_{b,v}$ is the probability that $b$’s realized value is $v$. This would result in a competitive ratio of $O(\gamma)$, for a scaling factor $\gamma$ large enough to absorb demand fluctuations with constant probability.
A key challenge is that such independent acceptance probabilities cannot be implemented directly by a static and anonymous menu, in which the mechanism must commit ex-ante to offering a fixed number of copies of each bundle at a given price. To do so, guided by the solution of EA-LP, we post no copies of bundle $S$ at prices below its important value $w_{S}$ and practically unlimited copies at the smallest price $w_{S}+1$ that is strictly larger than $S$’s important value. Thus, we rely on the scarcity of high-value buyers (as quantified by the $\gamma$-scaled down capacity constraints of EA-LP) to limit the number of item copies allocated to the buyers. The number of copies offered at $S$’s important value $w_{S}$ requires more care. Offering too many copies risks exhausting item capacities and blocking high-value buyers later in the sequence, while offering too few copies may result in a significant loss relative to the fractional benchmark. To balance these effects, we offer $\lfloor x_{S,w_{S}}\rfloor$ additional copies of a bundle $S$ at its important value $w_{S}$, where $x_{S,w_{S}}$ denotes the total EA-LP allocation to buyers demanding bundle $S$ at value $w_{S}$. If $x_{S,w_{S}}<1$, we randomize, with appropriate probability, between offering one additional copy of $S$ at price $w_{S}$ or not (see also Algorithm 1 in Section 6).
Our mechanism posts the static and anonymous menu described above, which consists of tuples $(S,w,c)$, denoting that $c$ copies of $S$ are available at price $w$. Let $w_{b}(S_{b})$ be the lowest price at which a copy of $S_{b}$ is available in the mechanism’s menu on the arrival of a buyer $b$. If $b$’s realized value $v_{b}$ is at least $w_{b}(S_{b})$, a corresponding menu entry is allocated to $b$ and the number of $S_{b}$’s copies available at price $w_{b}(S_{b})$ decreases by one. If all items $e\in S_{b}$ are still available (i.e., their residual capacity is positive), the bundle $S_{b}$ is allocated to $b$, the mechanism collects value $v_{b}$ and the residual capacity of all items $e\in S_{b}$ decreases by one.
We highlight that the number of times an item $e\in\mathcal{M}$ is included in bundles offered by the mechanism’s menu is typically much larger than $e$’s actual capacity. Hence, it may be the case that a buyer $b$ demands their bundle $S_{b}$ at a price currently available in the menu, but $b$ does not receive $S_{b}$ because some item $e\in S_{b}$ has been exhausted (i.e., $e$’s residual capacity is 0). To deal with this situation in the analysis, we introduce a (capacity) unconstrained version of our mechanism: a hypothetical process that completely ignores the item capacities and is only restricted by the mechanism’s menu entries (guided as above by the fractional optimal solution to $\gamma$-scaled EA-LP). The unconstrained mechanism is not hard to analyze and satisfies two key properties:
  1. 1.
    Its expected social welfare is within a constant factor from the optimal value of the $\gamma$-scaled EA-LP, denoted $\textnormal{{FracOpt}}_{\gamma}$. In turn, $\textnormal{{FracOpt}}_{\gamma}$ is within $O(\gamma)$ of the prophet’s expected social welfare.
  2. 2.
    For every item $e$, the number of $e$’s copies allocated to the buyers is at most $B/\gamma$ in expectation and (with an appropriate choice of $\gamma$) at most $B$ with constant probability.
The social welfare achieved by our mechanism equals the unconstrained welfare minus the expected value of blocked buyers, namely buyers receiving a menu entry in the unconstrained mechanism, but not their desired bundle due to the item capacity constraints. Due to Property (1), it suffices to upper bound the expected welfare lost due to blocked buyers by a constant fraction of $\textnormal{{FracOpt}}_{\gamma}$. To this end, we crucially exploit Property (2).
A key technical challenge in the analysis is that (due to the power of the almighty adversary) the availability of a buyer’s desired bundle depends, in a quite delicate way, on both the buyers’ realized values and the mechanism’s randomization on the number of a bundle’s $S$ copies posted at its important value $w_{S}$. Hence, we cannot deal with these two sources of randomness as independent anymore. Instead, we need to take some careful steps in order to separate the two sources of randomness. We believe that the careful analysis of our randomized menu construction might be of independent interest.